OpenStudy (anonymous):

solve using any method. 2x^2+x-5=0 preferably quadratic

OpenStudy (anonymous):

are you familiar with the quadratic formula?

OpenStudy (anonymous):

yeah i just want to make sure that I did it right and idk how to put in the sq root sign on here.

OpenStudy (anonymous):

\[\frac{ -1\pm \sqrt{1+4(2)(5)} }{ 2(2) }\]

OpenStudy (anonymous):

is that what your starting out formula looks like once you plug in a b and c?

OpenStudy (anonymous):

yeah except its minus 4 and negative 5

OpenStudy (anonymous):

\frac{ -1\pm \sqrt{1-4(2)(5)} }{ 2(2) }

OpenStudy (anonymous):

yeah my bad

OpenStudy (anonymous):

so you solved for the expression under the root?

OpenStudy (anonymous):

\[\frac{ -1\pm \sqrt{41} }{ 4 }\]

OpenStudy (anonymous):

yeah?

OpenStudy (anonymous):

yeah that is what I got thank you

OpenStudy (anonymous):

since you cannot simplify,